# Meson-nucleus bound states in quark meson coupling model

### v2

@@ II.2 Mesic-nuclei potentials
$V_{v}^{K^{-},D^{-},B^{-}}(r)=V_{\omega}^{q}(r)-\frac{1}{2}V_{\rho}^{q}(r)-A(r),$ (54)
$V_{v}^{\bar{K}^{0},\bar{D}^{0},\bar{B}^{0}}(r)=V_{\omega}^{q}(r)+\frac{1}{2}V_{\rho}^{q}(r),$ (55)
$V_{v}^{K^{0},D^{0},B^{0}}(r)=-(V_{\omega}^{q}(r)+\frac{1}{2}V_{\rho}^{q}(r))$ (56)
@@ III.3 $K(\bar{K})$, $D(\bar{D})$ and $B(\bar{B})$ mesic-nuclei potentials
Figure 4: Behavior of (a) the relative masses and (b)-(d) the vectorial interactions of $P^{-}$ ($K^{-}$, $D^{-}$, $B^{-}$), $\bar{P}^{0}$ ($\bar{K}^{0}$, $\bar{D}^{0}$, $\bar{B}^{0}$), and $P^{0}$ ($K^{0}$, $D^{0}$, $B^{0}$) mesons while captured inside the ${\rm{{}^{208}Pb}}$ nucleus. Results are presented for both the standard $\omega$ coupling ($V_{\omega}^{q}$) and the phenomenologically suggested $\omega$ coupling ($\tilde{V}_{\omega}^{q}=1.4^{2}\times V_{\omega}^{q}$).
@@ III.4 Mesic-nuclei bound states
$K^{-}(V_{\omega}^{q})$ $1s$7.50 -14.53 -23.03 -29.22 -29.81
$1p$$\times$ -4.98 -13.93 -23.10 -24.48
$K^{-}(V_{\omega}^{q},\;\rm{no\;Coulomb})$ $1s$2.22 -5.60 -5.68 -6.96 -6.41
$1p$$\times$ $\times$ $\times$ -2.51 -2.22
$K^{-}(\tilde{V}_{\omega}^{q})$ $1s$1.15 -2.54 -9.09 -10.93 -13.99
$1p$$\times$ -0.84 -8.12 -10.10 -13.00
$K^{-}(\tilde{V}_{\omega}^{q},\;\rm{no\;Coulomb})$ $1s$$\times$ $\times$ $\times$ $\times$ $\times$
@@
$\bar{K}^{0}(V_{\omega}^{q})$ $1s$2.93 -9.08 -16.07 -21.93 -32.11
$1p$$\times$ $\times$ -7.07 -14.62 -24.85
$\bar{K}^{0}(\tilde{V}_{\omega}^{q})$ $1s$$\times$ $\times$ $\times$ $\times$ -1.45
$1p$$\times$ $\times$ $\times$ $\times$ $\times$
$K^{0}(V_{\omega}^{q})$ $1s$75.56 -94.89 -94.74 -88.32 -76.21
$1p$37.35 -66.79 -75.72 -77.60 -66.24
$K^{0}(\tilde{V}_{\omega}^{q})$ $1s$119.16 -144.87 -140.51 -129.54 -111.73
$1p$74.96 -111.50 -$1$117.16 -99.95
Our results are limited to ground states, as higher states require consideration of spin-orbit interactions, which are beyond the scope of the current study. The results show the possible formation of the bound states for $K(\bar{K})$ (in Table III.4), $D(\bar{D})$ (in Table III.4) and $B(\bar{B})$ (in Table III.4) mesons with the ${\rm{{}^{16}O}}$, ${\rm{{}^{40}Ca}}$, ${\rm{{}^{90}Zr}}$, ${\rm{{}^{197}Au}}$, and ${\rm{{}^{208}Pb}}$ nucleus under the various conditions: (i) with the usual vector-isoscalar potential $V_{\omega}^{q}$, (ii) with phenomenologically driven vector-isoscalar potential $\tilde{V}_{\omega}^{q}$, (iii) with Coulomb interaction, and (iv) without Coulomb interaction. The Coulomb interaction provides more attractive potentials to the negatively charged mesons within the nucleus, contributing approximately 6 MeV, 9 MeV, 18 MeV, 23 MeV, and 23 MeV binding for the $1s$ state for $K^{-}$, $D^{-}$, and $B^{-}$ mesons with ${\rm{{}^{16}O}}$, ${\rm{{}^{40}Ca}}$, ${\rm{{}^{90}Zr}}$, ${\rm{{}^{197}Au}}$, and ${\rm{{}^{208}Pb}}$ nucleus, respectively.
@@
Due to the sole contribution of the Coulomb force, the system with the much more repulsive potential $\tilde{V}_{\omega}^{q}$ can also form bound states. This is due to the behavior of the wave functions for the potential $\tilde{V}_{\omega}^{q}$, which indicates that the negatively charged mesons are comparatively pushed out of the nucleus D1 and therefore the bound states are formed solely due to the Coulomb force. However, the neutral mesons are blind to the Coulomb force. Therefore, they only experience the nuclear forces, and the resulting bound states are more immersed within the nucleus, unlike the mesic atom states which orbit around the nucleus.
@@
$B^{-}(V_{\omega}^{q})$ $1s$21.15 -23.96 -31.38 -33.29 -33.31
$1p$17.89 -22.76 -29.11 -32.49 -32.73
$B^{-}(V_{\omega}^{q},\;\rm{no\;Coulomb})$ $1s$14.55 -14.21 -12.06 -10.47 -9.61
$1p$11.56 -13.18 -10.34 -10.14 -9.30
$B^{-}(\tilde{V}_{\omega}^{q})$ $1s$3.28 -5.38 -11.42 -13.74 -17.58
$1p$3.01 -5.12 -11.28 -13.64 -17.49
$B^{-}(\tilde{V}_{\omega}^{q},\;\rm{no\;Coulomb})$ $1s$$\times$ $\times$ $\times$ $\times$ $\times$
@@
$\bar{B}^{0}(V_{\omega}^{q})$ $1s$16.53 -21.21 -27.47 -26.24 -37.24
$1p$12.95 -18.37 -22.74 -25.80 -36.49
$\bar{B}^{0}(\tilde{V}_{\omega}^{q})$ $1s$$\times$ $\times$ $\times$ $\times$ -3.87
$1p$$\times$ $\times$ $\times$ $\times$ -3.53
$B^{0}(V_{\omega}^{q})$ $1s$113.02 -123.48 -112.35 -96.58 -84.43
$1p$106.96 -113.61 -109.22 -94.59 -81.92
$B^{0}(\tilde{V}_{\omega}^{q})$ $1s$162.08 -182.06 -161.45 -139.61 -122.42
$1p$155.93 -167.80 -158.23 -137.06 -118.95
@@
With the usual $\omega$ coupling, the vector-isovector interaction produces a strong attractive potential for $\bar{K}^{0}$, $\bar{D}^{0}$, and $\bar{B}^{0}$ mesons. Figure 6, shows the obtained $1s$ and $1p$ states’ wave functions for these mesons under the usual vector-isoscalar potential $V_{\omega}^{q}$. To understand the contribution of vector-isovector interaction in terms of binding energies, one can compare the results with the negatively charged mesons without the Coulomb interaction, where the vector-isovector interaction is repulsive. It shows approximately 1 MeV, 6 MeV, 10 MeV, 16 MeV, and 26 MeV bindings are induced for the $\bar{K}^{0}$, $\bar{D}^{0}$, and $\bar{B}^{0}$ mesons while captured inside ${\rm{{}^{16}O}}$, ${\rm{{}^{40}Ca}}$, ${\rm{{}^{90}Zr}}$, ${\rm{{}^{197}Au}}$, and ${\rm{{}^{208}Pb}}$ nucleus, respectively. However, with the repulsive $\omega$ interaction, these mesons fail to form bound states unless the nucleus is heavy (e.g., ${\rm{{}^{208}Pb}}$), where they tend to form lightly bound states. In comparison to the results presented in Ref. D1 , the observed small binding energies are attributed to the inclusion of scalar-isovector potential. Essentially the potential $V_{\delta}$ induces opposite interactions between the mesons within the isospin doublets of $K(\bar{K})$, $D(\bar{D})$, and $B(\bar{B})$ mesons. As a consequence, it modifies the binding energies depending on the nature of their interactions. It is important to note that if the obtained states have small binding energies, they are more likely to mix with the continuum energy spectrum if their corresponding half-widths are greater laura1 . Therefore, this leads to indistinct states on the energy spectrum. Furthermore, the $K^{0}$, $D^{0}$, and $B^{0}$ mesons form deeply bound states with the nuclei. With $\tilde{V}_{\omega}^{q}$, the bound-state energies are even more compared to those calculated with $V_{\omega}^{q}$. However, as indicated in Ref. D1 , due to the strong absorption of these states, the anticipated large width could make experimental observation difficult. The opposite vector-isoscalar ($\omega$) interaction for $K^{0}\;(D^{0},B^{0})$ and $\bar{K}^{0}\;(\bar{D}^{0},\bar{B}^{0})$ leads to a significant difference in their binding energies, reflecting the presence of a strong Lorentz vector mean field. The results show the possible formation of bound states is more permissible with the heavier mesons, even when the potentials are comparable to each other. This represents the scalar interaction strengths of these mesons within the nuclei. Owing to the constituents of the open beauty mesons, they can form more strongly bound states with the nucleus. As depicted by the last column of Fig. 6, for these states, the meson is much closer to the nuclei, which allows these states to probe even smaller variations in the nuclear matter properties krein1 .

### v3

@@ II.2 Mesic-nuclei potentials
$V_{v}^{K^{-},B^{-}}(r)$ $=-V_{\omega}^{q}(r)-\tfrac{1}{2}V_{\rho}^{q}(r)-A(r),$ $V_{v}^{D^{-}}(r)$ $=V_{\omega}^{q}(r)-\tfrac{1}{2}V_{\rho}^{q}(r)-A(r),$ (54)
$V_{v}^{\bar{K}^{0},\bar{B}^{0}}(r)$ $=-V_{\omega}^{q}(r)+\tfrac{1}{2}V_{\rho}^{q}(r),$ $V_{v}^{\bar{D}^{0}}(r)$ $=V_{\omega}^{q}(r)+\tfrac{1}{2}V_{\rho}^{q}(r),$ (55)
$V_{v}^{K^{0},B^{0}}(r)$ $=V_{\omega}^{q}(r)-\tfrac{1}{2}V_{\rho}^{q}(r),$ $V_{v}^{D^{0}}(r)$ $=-V_{\omega}^{q}(r)-\tfrac{1}{2}V_{\rho}^{q}(r).$ (56)
@@ III.3 $K(\bar{K})$, $D(\bar{D})$ and $B(\bar{B})$ mesic-nuclei potentials
Figure 4: Behavior of (a) the relative masses and (b)-(d) the vectorial interactions of $P^{-}$ ($K^{-}$, $D^{-}$, $B^{-}$), $\bar{P}^{0}$ ($\bar{K}^{0}$, $\bar{D}^{0}$, $\bar{B}^{0}$), and $P^{0}$ ($K^{0}$, $D^{0}$, $B^{0}$) mesons while captured inside the ${\rm{{}^{16}O}}$ nucleus. Results are presented for both the standard $\omega$ coupling ($V_{\omega}^{q}$) and the phenomenologically suggested $\omega$ coupling ($\tilde{V}_{\omega}^{q}=1.4^{2}\times V_{\omega}^{q}$).
@@ III.4 Mesic-nuclei bound states
$K^{-}(V_{\omega}^{q})$ $1s$81.67 -105.17 -112.39 -112.28 -100.75
$1p$42.95 -76.38 -92.81 -100.58 -90.02
$K^{-}(V_{\omega}^{q},\;\rm{no\;Coulomb})$ $1s$74.98 -94.36 -93.29 -87.91 -75.81
$1p$36.66 -66.16 -74.40 -77.14 -65.79
$K^{-}(\tilde{V}_{\omega}^{q})$ $1s$125.34 -155.29 -158.45 -153.75 -136.48
$1p$80.67 -121.28 -136.56 -140.46 -123.96
$K^{-}(\tilde{V}_{\omega}^{q},\;\rm{no\;Coulomb})$ $1s$118.54 -144.29 -139.18 -129.12 -111.32
$1p$74.17 -110.80 -117.89 -116.68 -99.49
$\bar{K}^{0}(V_{\omega}^{q})$ $1s$77.44 -104.52 -106.16 -106.30 -104.72
$1p$38.48 -73.37 -88.31 -93.84 -93.62
$\bar{K}^{0}(\tilde{V}_{\omega}^{q})$ $1s$121.25 -155.34 -151.82 -147.99 -140.59
$1p$76.43 -119.42 -133.62 -133.97 -127.84
$K^{0}(V_{\omega}^{q})$ $1s$2.52 -5.93 -5.37 -7.29 -6.74
$1p$$\times$ $\times$ $\times$ -2.80 -2.55
$K^{0}(\tilde{V}_{\omega}^{q})$ $1s$$\times$ $\times$ $\times$ $\times$ $\times$
@@
Our results are limited to ground states, as higher states require consideration of spin-orbit interactions, which are beyond the scope of the current study. The results show the possible formation of the bound states for $K(\bar{K})$ (in Table III.4), $D(\bar{D})$ (in Table III.4) and $B(\bar{B})$ (in Table III.4) mesons with the ${\rm{{}^{16}O}}$, ${\rm{{}^{40}Ca}}$, ${\rm{{}^{90}Zr}}$, ${\rm{{}^{197}Au}}$, and ${\rm{{}^{208}Pb}}$ nucleus under the various conditions: (i) with the usual vector-isoscalar potential $V_{\omega}^{q}$, (ii) with phenomenologically driven vector-isoscalar potential $\tilde{V}_{\omega}^{q}$, (iii) with Coulomb interaction, and (iv) without Coulomb interaction. The Coulomb interaction provides more attractive potentials to the negatively charged mesons within the nucleus, contributing approximately 6-7 MeV, 9-11 MeV, 18-19 MeV, 23-25 MeV, and 24-26 MeV binding for the $1s$ state for $K^{-}$, $D^{-}$, and $B^{-}$ mesons with ${\rm{{}^{16}O}}$, ${\rm{{}^{40}Ca}}$, ${\rm{{}^{90}Zr}}$, ${\rm{{}^{197}Au}}$, and ${\rm{{}^{208}Pb}}$ nucleus, respectively.
@@
$B^{-}(V_{\omega}^{q})$ $1s$119.77 -134.88 -130.71 -121.83 -110.11
$1p$113.62 -124.60 -127.52 -119.39 -107.22
$B^{-}(V_{\omega}^{q},\;\rm{no\;Coulomb})$ $1s$112.73 -123.16 -110.82 -96.31 -84.16
$1p$106.67 -113.31 -101.35 -94.81 -81.65
$B^{-}(\tilde{V}_{\omega}^{q})$ $1s$168.84 -193.55 -179.95 -165.04 -148.22
$1p$162.63 -179.01 -176.69 -162.06 -144.43
$B^{-}(\tilde{V}_{\omega}^{q},\;\rm{no\;Coulomb})$ $1s$161.79 -181.74 -160.01 -139.34 -122.15
$1p$155.64 -167.49 -156.81 -136.80 -118.68
$\bar{B}^{0}(V_{\omega}^{q})$ $1s$115.66 -135.22 -122.03 -117.32 -114.61
$1p$109.45 -125.92 -118.73 -114.22 -111.59
$\bar{B}^{0}(\tilde{V}_{\omega}^{q})$ $1s$164.77 -193.33 -$1$160.74 -152.78
$1p$158.51 -180.32 -167.67 -157.27 -149.15
$B^{0}(V_{\omega}^{q})$ $1s$14.80 -14.43 -10.98 -10.70 -9.84
$1p$11.79 -13.40 -9.44 -10.37 -9.52
$B^{0}(\tilde{V}_{\omega}^{q})$ $1s$$\times$ $\times$ $\times$ $\times$ $\times$
@@
Due to the sole contribution of the Coulomb force, the $D^{-}$ meson with the much more repulsive potential $\tilde{V}_{\omega}^{q}$ can also form bound states. This is due to the behavior of the wave functions for the potential $\tilde{V}_{\omega}^{q}$, which indicates that the negatively charged $D^{-}$ mesons are comparatively pushed out of the nucleus and therefore the bound states are formed solely due to the Coulomb force. With the same light quark structure, ${K}^{0}$ and ${B}^{0}$ mesons, since are blind to the Coulomb force, are unable to form the bound states for $\tilde{V}_{\omega}^{q}$. Therefore, the neutral states only experience the nuclear forces, and the resulting bound states are more immersed within the nucleus. With the usual $\omega$ coupling, the vector-isovector interaction produces a strong attractive potential for the $\bar{K}^{0}$ and $\bar{B}^{0}$ mesons compared to $\bar{D}^{0}$. The resulting binding energies become higher for the $\bar{K}^{0}$ and $\bar{B}^{0}$ mesons for a more repulsive $\omega$ coupling ($\tilde{V}_{\omega}^{q}$). However, for such scenario, $\bar{D}^{0}$ is less likely to bound within nuclei. It shows approximately 1 MeV, 6 MeV, 10 MeV, 16 MeV, and 26 MeV bindings are induced for the $\bar{D}^{0}$ meson while captured inside ${\rm{{}^{16}O}}$, ${\rm{{}^{40}Ca}}$, ${\rm{{}^{90}Zr}}$, ${\rm{{}^{197}Au}}$, and ${\rm{{}^{208}Pb}}$ nucleus, respectively, while for $\bar{K}^{0}$ and $\bar{B}^{0}$ mesons the obtained binding energies are quite large. The values for $\bar{K}^{0}$ and $\bar{B}^{0}$ can be seen from Tables III.4 and III.4, respectively. Notably, with the repulsive $\omega$ interaction, a few scenarios, which fail to form bound states, tend to form very light bound states with the heavy nuclei (e.g., ${\rm{{}^{208}Pb}}$). In comparison to the results presented in Ref. D1 , the observed small binding energies are attributed to the inclusion of scalar-isovector potential. Essentially, the potential $V_{\delta}$ induces opposite interactions between the mesons within the isospin doublets of $K(\bar{K})$, $D(\bar{D})$, and $B(\bar{B})$ mesons. As a consequence, it modifies the binding energies depending on the nature of their interactions. It is important to note that if the obtained states have small binding energies, they are more likely to mix with the continuum energy spectrum if their corresponding half-widths are greater laura1 . Therefore, this leads to indistinct states on the energy spectrum. Furthermore, the $K^{-}$, $D^{0}$, and $B^{-}$ mesons form deeply bound states with the nuclei. With $\tilde{V}_{\omega}^{q}$, the bound-state energies are even more compared to those calculated with $V_{\omega}^{q}$. However, as indicated in Ref. D1 , due to the strong absorption of these states, the anticipated large width could make experimental observation difficult. The opposite vector-isoscalar ($\omega$) interaction for $K^{0}\;(D^{0},B^{0})$ and $\bar{K}^{0}\;(\bar{D}^{0},\bar{B}^{0})$ leads to a significant difference in their binding energies, reflecting the presence of a strong Lorentz vector mean field. The results show the possible formation of bound states is more permissible with the heavier mesons, even when the potentials are comparable to each other. This represents the scalar interaction strengths of these mesons within the nuclei. Owing to the constituents of the open beauty mesons, they can form more strongly bound states with the nucleus. As depicted by the last column of Fig. 6, for these states, the meson is much closer to the nuclei, which allows these states to probe even smaller variations in the nuclear matter properties krein1 .